On January 1, 2018, Instaform, Inc., issued 10% bonds with a face amount of $53 million, dated January 1. The bonds mature in 2037 (20 years). The market yield for bonds of similar risk and maturity is 12%. Interest is paid semiannually. (FV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided.)

Required:
1-a.
Determine the price of the bonds at January 1, 2018.
1-b. Prepare the journal entry to record their issuance by Instaform.
2-a. Assume the market rate was 9%. Determine the price of the bonds at January 1, 2018.
2-b. Assume the market rate was 9%. Prepare the journal entry to record their issuance by Instaform.
3. Assume Broadcourt Electronics purchased the entire issue in a private placement of the bonds. Using the data in requirement 2, prepare the journal entry to record the purchase by Broadcourt.
Answer #1

1-a. Determine the price of the bonds at January 1, 2018.

Present value of interest= $53000000*10%*6/12*15.04630= $39872695

Present value of Principal= $53000000*0.09722= $5152660

Present value of bonds= Present value of Principal+Present value of interest

= $5152660+39872695= $45025355

1-b. Prepare the journal entry to record their issuance by Instaform.

Date Account titles and explanation Debit Credit
January 1, 2018 Cash $45025355
Discount on bonds payable (53000000-45025355) $7974645
Bonds payable $53000000
(To record issuance of bonds)

2-a. Assume the market rate was 9%. Determine the price of the bonds at January 1, 2018.

Present value of interest= $53000000*10%*6/12*18.40158= $48764187

Present value of Principal= $53000000*0.17193= $9112290

Present value of bonds= Present value of Principal+Present value of interest

= $9112290+48764187= $57876477

2-b. Assume the market rate was 9%. Prepare the journal entry to record their issuance by Instaform.

Date Account titles and explanation Debit Credit
January 1, 2018 Cash $57876477
Premium on bonds payable (57876477-53000000) $4876477
Bonds payable $53000000
(To record issuance of bonds)

3. Assume Broadcourt Electronics purchased the entire issue in a private placement of the bonds. Using the data in requirement 2, prepare the journal entry to record the purchase by Broadcourt.

Date Account titles and explanation Debit Credit
January 1, 2018 Investment in bonds $53000000
Premium on bonds (57876477-53000000) $4876477
Cash $57876477
(To record purchase of bonds)

Present value of interest= $53000000*10%*6/12*18.40158= $48764187

Present value of Principal= $53000000*0.17193= $9112290

Present value of bonds= Present value of Principal+Present value of interest

= $9112290+48764187= $57876477